OpenStudy (anonymous):

Trigonometry Help Prove the Identity: cos^2 x + 1 = 2cos^2 x + sin^2 x Any help is greatly appreciated with this one.

OpenStudy (anonymous):

do u know cos^2 x + sin^2 x =1

OpenStudy (anonymous):

yeah i just saw that so on the right it beceomes =2(1) which =2? or is that wrong?

OpenStudy (anonymous):

cos^2 x + 1 = cos^2 x + cos^2 x + sin^2 x = 2cos^2 x + sin^2 x

OpenStudy (anonymous):

well im totally wrong :\

OpenStudy (akashdeepdeb):

Did you understand it NOW? :D

OpenStudy (anonymous):

mattricked the lasts message you sent, was that working on the left or the right side?

OpenStudy (anonymous):

ahh ok nevermind, i see it now. thanks guys

OpenStudy (anonymous):

yw